OpenStudy (anonymous):

Use the unit circle to find the inverse function value in degrees. cos -1((sqrt)3/2) A. 60° B. 30° C. 240° D. 150°

OpenStudy (anonymous):

can someone just show me how to get my answer without telling it to me?

OpenStudy (anonymous):

@satellite73 @zepdrix

OpenStudy (anonymous):

@Loser66

zepdrix (zepdrix):

\[\Large\bf\sf \cos330^o\quad=\quad \cos30^o\quad=\quad \frac{\sqrt3}{2}\]

zepdrix (zepdrix):

The inverse cosine function, in order to be a function, has it's domain restricted from 0 to pi. So the inverse cosine can only produce angles in the first 2 quadrants. Which angle do we want? 330 or 30 degrees?

OpenStudy (anonymous):

30 degrees?

zepdrix (zepdrix):

Woops, domain restricted from -1 to 1 I think I meant to say :O Err blah it's not too important to what we're doing...

zepdrix (zepdrix):

Yes, because 30 degrees is in the first quadrant. 330 degrees is in the 4th quadrant, which is not in the range of our function arccosine, which only gives us angles in the first 2 quadrants.

OpenStudy (anonymous):

Ok so now what do we do?

zepdrix (zepdrix):

We choose B silly! :3

OpenStudy (anonymous):

*facepaw* Oh XDD

OpenStudy (anonymous):

Sorry i didn't pay attention to my answers i put

zepdrix (zepdrix):

heh